OpenStudy (anonymous):

Please help and explain.. Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x^2 + 13x + 15 2x^2 + 27x + 13 5x^2 − 2x − 7 10x^2 − 7x − 3 8x^2 − 11x + 3

OpenStudy (acacia):

you need to find numbers that multiply to 2 and 15 and whose products add up to thirteen,etc. 2x^2 +13x +15 = (2x +3)(x +5) 2x^2 +27x +13 = (2x +1)(x +13) 5x^2 -2x -7 = (5x -7)(x +1) 10x^2 -7x -3 = (x -1)(10x +3) 8x^2 -11x +3 = (x -1)(8x -3)

OpenStudy (anonymous):

Would it be 2x^2 + 27x + 13 && 10x^2 − 7x − 3 ?

OpenStudy (acacia):

2x^2 + 13x + 15 1.multiply 2*15 to get: x^2+13x+30 2. factor: [x+10][x+3]=0 where the 2#'s must =+13 when added and +30 when multiplied. 3. put back the 2 to get: [2x+10][2x+3]=0 4. reduce if you can: [2x+10] is the same as [x+5]; [2x+3] stays as is. ***^^^put that into sentences^^^*** 2x^2 + 27x + 13 x^2+27x+26 [x+26][x+1] [2x+26][2x+1] [x+13][2x+1] answer 5x^2 − 2x − 7 x^2-2x-35 [x-7][x+5] [5x-7][5x+5] [5x-7][x+1] answer 10x^2 − 7x − 3 x^2-7x-30 [x-10][x+3] [10x-10][10x-3] [x-1][10x-3] answer 8x^2 − 11x + 3 x^2-11x+24 [x-8][x-3] [8x-8][8x-3] [x-1][8x-3] answer i used the same method for all of them.